Tickets to a local movie were sold at $4 for adults and $2.50 for students. If 59 tickets were sold for a total of $222.50, how many student tickets were sold?
Found 2 solutions by mananth, robertb:
Answer by mananth(16949)

Tickets to a local movie were sold at $4 for adults and $2.50 for students. If 59 tickets were sold for a total of $222.50, how many student tickets were sold?
$4 for adults
$2.50 for students.
59 tickets were sold
total of $222.50,
adults x
students y
x+y=59.............1
4x+2.5y=222.5......2
...
multiply (1) by -4
-4x-4y=-236
add to (2)
-4x+4x-4y+2.5y=222.5-236
-1.5y=-13.5
/1.5
y=9 students tickets
adult tickets = 50

Answer by robertb(5830)

4x + 2.50(59 - x) = 222.50
4x + 147.50 - 2.50x = 222.50
1.50x = 75
x = 50, the number of adult tickets sold, and
59 - x = 9, the number of student tickets sold.